OpenStudy (anonymous):

2x(x+3xy) =2x+x+2x+3xy =2x^2+3xy^2 2x^2+3xy^2 Did I get this right?

hartnn (hartnn):

\(2x \times x + 2x \times 3xy \\ = 2x^2 + (2\times 3) x \times x \times y \\ =2x^2 +6x^2 y\)